# Benign Overfitting for $\alpha$ Sub-exponential Input

### v1

@@ 5 Conclusion
    Our results demonstrate that as the number of dimensions $p$ increases and the feature vector magnitude $\|\mu\|$ scales as $\Theta(p^{d})$ for $d\in(1/4,1/2]$, the misclassification error approaches the noise level $\eta$.
@@ Acknowledgements
Appendix A Proof of lemmas
A.1 Concentration inequality
@@
A.2 Proof of Lemma 10
In this section, we prove Lemma 10 by using concentration inequalities from Section A.1. We assume Assumptions (A1)-(A4) hold, and decompose Lemma 10 into six different parts. We prove that each separate lemma holds with probability at least $61-\delta/$ 6.
@@ A.2 Proof of Lemma 10
$2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right)=2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right)$
@@
Combining these inequalities, we obtain:
@@
$=\frac{\kappa p}{4}-\frac{p}{C}$
@@
$\|z_{k}\|^{2}>\frac{\kappa p}{4}-\frac{\kappa p}{8}=\frac{\kappa p}{8}$ (3)
@@
$<\left(\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)p$
@@
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)$
@@
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq t\right]\leq 2n^{2}\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}]$
@@
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{24}+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}]$
@@
$\leq 3\|\mu\|^{2}+c\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$
@@
$=\frac{1}{c}\left(p-2c^{2}n\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)$
By Assumption (A3) and (A4), for large enough $C$ we have

### v2

@@ 6 Conclusion
    Our results demonstrate that as the number of dimensions $p$ increases and the feature vector magnitude $\|\mu\|$ scales as $\Theta(p^{d})$ for $d\in(1/4,1/2]$, the misclassification error approaches the noise level $\eta$ even under the heavy-tailed setting.
@@
  4. 4.
    We showed that the upper bound on the learning rate for benign overfitting, and demonstrated that when $n$ is fixed, the bound is of order $p^{-1}$, while in the case where $n,p,\|\mu\|$ are large, we observed that the upper bound decreases as $\alpha$ increases.
  5. 5.
    By conducting simulations, we confirmed that the relationship between the number of parameters, the tail heaviness, and the learning rate when benign overfitting occurs follows the same trend as that derived theoretically.
@@ Acknowledgements
Supplementary Materials
Appendix A Concentration inequality
@@
Appendix B Missing proofs
B.1 Proofs of lemmas used in the proof of Theorem 4
B.1.1 Proof of Lemma 12
Proof of Lemma 12.
By definition of $\tilde{z}_{k}$ and $z_{k}$,
Together with the definition of $\xi_{k}$,
B.1.2 Proof of Lemma 13
Proof of Lemma 13.
Following the proof of Chatterji and Long [7], we have
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right].$
Applying Proposition 18, there exists a positive constant $c$ such that
which completes our proof. ∎
B.1.3 Proof of Lemma 14
In this section, we prove Lemma 14 by using concentration inequalities from Section A. We assume assumptions (A1)-(A4) hold, and decompose Lemma 14 into six different parts. We prove that each separate lemma holds with probability at least $61-\delta/$ 6.
@@ B.1.3 Proof of Lemma 14
$2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right)=2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right).$
@@
Combining these inequalities, we obtain
@@
$=\frac{\kappa p}{4}-\frac{p}{C}.$
@@
$\|z_{k}\|^{2}>\frac{\kappa p}{4}-\frac{\kappa p}{8}=\frac{\kappa p}{8}.$ (2)
@@
$<\left(\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)p.$
@@
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right),$
@@
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq t\right]\leq 2n^{2}\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}].$
@@
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{24}+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}].$
@@
$\leq 3\|\mu\|^{2}+c\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}.$
@@
$=\frac{1}{c}\left(p-2c^{2}n\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right).$
By assumptions (A3) and (A4), for large enough $C$ we have